OpenStudy (anonymous):

Simplify. Express the quotient as a radical expression. ninth root of x to the seventh power over ninth root of x to the fifth power ninth root of x squared ninth root of x to the twelfth power zero root of x squared eighty first root of x to the thirty fifth power do i subtract or multiply?????

OpenStudy (anonymous):

i don't understand... subtract or multiply what? are these 5 separate problems?

OpenStudy (anonymous):

no the bottom is the answer list

OpenStudy (anonymous):

cus if i subtract 7-5 it would =2 wich is an answer but if i multiply then it would be 35 which is also an answer

OpenStudy (anonymous):

oh i see.... this is the problem: "ninth root of x to the seventh power over ninth root of x to the fifth power" and the following 4 lines are the answer choices...???

OpenStudy (anonymous):

\[\sqrt[9]{x7} \over \sqrt[9]{x5}\]

OpenStudy (anonymous):

yea exactly

OpenStudy (anonymous):

writing the top and bottom as powers with rational exponents, you'll get: \(\huge \frac{x^{\frac{7}{9}}}{x^{\frac{5}{9}}} \).... subtract the exponents... to get \(\large x^{\frac{2}{9}} \) now write that out in radical form

OpenStudy (anonymous):

ohhhh snap man thanks i got it!!!!!!!! :D